TheKayla:

Solve the algebra equation for x and y: x - 3y - 3 = 0 3x - 9y - 2 = 0

TheKayla:

@Elsa help???

TheKayla:

@Elsa213

TheKayla:

Assistance?

enbyinyourbasement:

I used the substitution method and, unless I missed something, I got no possible answers.

jhonyy9:

@enbyinyourbasement wrote:
I used the substitution method and, unless I missed something, I got no possible answers.
ok I believe you but please post here your work ,calcules ! thank you !

jhonyy9:

x -3y -3 = 0 add to both sides 3 and will get x -3y = 3 3x -9y -2 = 0 add to both sides 2 and will get 3x -9y = 2 x -3y = 3 3x -9y = 2 and now solve this like a system of two equations

surjithayer:

x-3y=3 multiply by 3 3x-9y=9 also 3x-9y=2 so system is inconsistent. Hence no solution.